# Quantum Corrections to $\eta/s$ from JT Gravity

### v2

@@ 1 Introduction
Thus, our logic in this paper is the following. We compute quantum corrections to the IR retarded Green’s function $\mathcal{G}_{R}(\omega,T)$, working in two different temperature regimes – one which we refer to as semiclassical, for which the temperature scale dominates over the quantum scale, and one which we refer to as quantum, in which the quantum effects dominate. We should note that, despite being in a low temperature regime, we always work with $\omega\ll T$ to ensure that hydrodynamics is well defined. We then relate $\mathcal{G}_{R}(\omega,T)$ to the retarded Green’s function $G_{R}(\omega,T)$ in the UV, using Faulkner:2009wj ; Faulkner:2013bna ; Faulkner:2010da , and use Kubo’s formula to extract the shear viscosity. We should stress that – unlike in the semiclassical regime, which is well-understood – our computation of $\eta$ in the deep quantum regime should be interpreted with caution, since a new framework is likely needed to accurately capture hydrodynamics. In this paper we will not address this challenge, but rather present our computations in both regimes – hoping that the results in the quantum regime might also provide guidance.
@@
As we will see, the quantum corrections will generate a non-trivial temperature dependence for $\eta/s$, unlike the tree-level case for which $\eta/s=1/4\pi$. Moreover, we will find that $\eta/s$ reaches a minimum which violates the KSS bound in the semi-classical regime (and not in the deep quantum regime), suggesting that even a small amount of quantum corrections might lead to a lower bound. We will show that the minimum can be traced to a maximum in the quantum corrected entropy, while the viscosity itself is monotonic. We should also mention that recent studies PandoZayas:2025snm ; Nian:2025oei have investigated quantum corrections to hydrodynamic transport coefficients arising from the JT mode at low temperatures [^4] The earlier work Jensen:2016pah also attempted to rewrite a generic theory of gravity near the AdS2 throat as a novel hydrodynamics coupled to the correlation functions of a conformal quantum mechanics.. As we show explicitly in Section 4, our results for $\eta/s$ are in agreement with those of PandoZayas:2025snm .
@@ 4.2 Quantum corrected retarded Green’s function
We will analyze the Green’s function separately in two distinct temperature regimes. First, we will examine the semi-classical limit $T\gg\frac{1}{C}$ in which the quantum effects are much smaller than the temperature scale, and then the opposite quantum regime $T\ll\frac{1}{C}$ in which quantum effects dominate over temperature. However, one should keep in mind that the results obtained in the quantum regime should be interpreted with caution, as a new theoretical framework may be needed for a proper description of hydrodynamics and of the shear viscosity (we return to this point in the Conclusions). Nonetheless, we present our $\eta/s$ results for $TC\ll 1$ since they may still offer useful physical intuition, with the understanding that a more thorough analysis could lead to a different behavior in this particular temperature range. Throughout the analysis we focus on the regime $\omega\ll M$, which implies [^11] One can show that $M\propto T^{2}$ and $M\propto T$ at the saddle points, in the semiclassical and quantum regimes respectively. $\omega\ll T$, ensuring that the frequency remains the lowest energy scale in the system and that hydrodynamics is well-defined (subject to the caveat above).
@@
    We evaluate this integral using a saddle-point approximation and Fourier transformation to obtain the quantum-corrected retarded Green’s function,
    $\hskip-28.45274pt\langle\mathcal{G}_{R}(\omega,\beta)\rangle\approx-(2\pi)^{2\ell-1}\left[2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]^{1-2\ell}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right)}{\Gamma\left(1-\ell-i\frac{\omega}{2\pi}2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right)}\,.$ (4.17)
@@
    We evaluate this integral also using a saddle-point approximation and Fourier transformation to obtain the quantum-corrected retarded Green’s function, with the final result given by:
    $\langle\mathcal{G}_{R}(\omega,\beta)\rangle=-\frac{(\ell-1)!^{2}}{(2C)^{2\ell-1}}\frac{8}{\sqrt{\pi}e}\left(\frac{\beta}{2C}\right)^{\frac{1}{2}}\frac{\Gamma\left(\ell-i\left(\frac{1}{2}2C\omega\sqrt{\frac{\beta}{2C}}\right)\right)}{\Gamma\left(1-\ell-i\left(\frac{1}{2}2C\omega\sqrt{\frac{\beta}{2C}}\right)\right)}\,.$ (4.20)
    Details of the derivation are described in Appendix A.2.
@@ 4.2.1 A quantum corrected $\ell^{\prime}$
    $\hskip-28.45274pt\langle\mathcal{G}_{R}(\omega,\beta)\rangle\approx-(2\pi)^{2\ell-1}\left[2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]^{1-2\ell}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right)}{\Gamma\left(1-\ell-i\frac{\omega}{2\pi}2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right)}\,.$ (4.22)
@@
    $F_{1}(\beta,\ell,\omega)$ $=-\frac{\Gamma(\frac{3}{2}-\ell)}{\Gamma(\frac{1}{2}+\ell)}(\pi)^{2\ell-1}(2C)^{1-2\ell}\frac{\Gamma(\ell-\frac{i\omega}{2\pi T})}{\Gamma(1-\ell-\frac{i\omega}{2\pi T})}\,$, (4.24)
    $F_{2}(\beta,\ell,\omega)$ $=-(2\pi)^{2\ell-1}(2C)^{1-2\ell}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right)}{\Gamma\left(1-\ell-i\frac{\omega}{2\pi}2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right)}\,.$ (4.25)
@@
    $\ell_{1}^{\prime}=\frac{1}{2}-\frac{{\rm ln}\left(\frac{F_{2}(\beta,\ell,\omega)}{F_{1}(\beta,\ell,\omega)}\right)+(1-2\ell)\,{\rm ln}\left[\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]}{2\,{\rm ln}\,\frac{\beta}{2C}}\,$, (4.27)
@@
    $\langle\mathcal{G}_{R}(\omega,\beta)\rangle=-\frac{(\ell-1)!^{2}}{(2C)^{2\ell-1}}\frac{8}{\sqrt{\pi}e}\left(\frac{\beta}{2C}\right)^{\frac{1}{2}}\frac{\Gamma\left(\ell-i\left(\frac{1}{2}2C\omega\sqrt{\frac{\beta}{2C}}\right)\right)}{\Gamma\left(1-\ell-i\left(\frac{1}{2}2C\omega\sqrt{\frac{\beta}{2C}}\right)\right)}\,.$ (4.28)
@@
    $\langle\mathcal{G}_{R}(\omega,\beta)\rangle$ $=F_{3}(\beta,\ell,\omega)\times\left(\frac{\beta}{2C}\right)^{\frac{1}{2}}=F_{1}(\beta,\ell,\omega)\times\left(\frac{\beta}{2C}\right)^{1-2\ell_{2}^{\prime}}\,$, (4.29)
@@
    $F_{3}(\beta,\ell,\omega)$ $=-\frac{(\ell-1)!^{2}}{(2C)^{2\ell-1}}\frac{8}{\sqrt{\pi}e}\frac{\Gamma\left(\ell-i\left(\frac{1}{2}2C\omega\sqrt{\frac{\beta}{2C}}\right)\right)}{\Gamma\left(1-\ell-i\left(\frac{1}{2}2C\omega\sqrt{\frac{\beta}{2C}}\right)\right)}\,.$ (4.30)
@@
    $\ell_{2}^{\prime}=\frac{1}{4}+\frac{1}{2}\frac{\log{\left(\frac{F_{1}(\beta,\ell,\omega)}{F_{3}(\beta,\ell,\omega)}\right)}}{\log{\left(\frac{\beta}{2C}\right)}}\,.$ (4.31)
@@ 4.3 Quantum corrected shear viscosity and entropy density
To better illustrate the behavior of the shear viscosity, we write down its approximated expression at high temperature $T\gg\frac{1}{C}$,
@@
where $\cdots$ denotes $\mathcal{O}(r_{0}^{-2})$ and $\mathcal{O}((CT)^{-2})$ terms. We note that this expression (4.36) for the quantum-corrected $\eta^{\prime}$ in the semiclassical region is consistent with that of the recent complementary work PandoZayas:2025snm . However, the full expression (4.3) is generically more involved, and captures additional effects to those that appear to leading order in the expansion (4.36).
On the other hand, in the quantum regime $T\ll\frac{1}{C}$ we have
$\eta^{\prime}=\frac{1}{2\kappa^{2}}\left(\frac{r_{h}}{L}\right)^{2}T^{2\ell_{2}^{\prime}-2}$ $=\frac{1}{2\kappa^{2}}\left(\frac{r_{h}}{L}\right)^{2}\times\frac{(\ell-1)!^{2}}{(2C)^{2\ell-2}}\frac{2}{\sqrt{\pi}e}(CT)^{-1}$ (4.37)
$\approx\frac{1}{2\kappa^{2}}\left(\frac{r_{0}}{L}\right)^{2}\times\frac{(\ell-1)!^{2}}{(2C)^{2\ell-2}}\frac{2}{\sqrt{\pi}e}(CT)^{-1}+\cdots\,$, (4.38)
where $\cdots$ denotes higher orders of $\mathcal{O}(r_{0}^{-1})$. We will come back to these expressions for the shear viscosity in the next section, where we make a direct comparison to the behavior of the absorption cross-section. Also, we should stress that when we plot the viscosity to entropy ratio we will use the full expression (4.3) for the shear viscosity, and not the approximations above. In Figure 2 we plot the ratio of the quantum-corrected shear viscosity $\eta^{\prime}$ to the classical shear viscosity $\eta$ in the two distinct temperature regimes.
@@
Figure 5: The left panel shows the temperature dependence of the quantum-corrected shear viscosity-entropy ratio in the quantum regime $CT\ll 1$. The dashed blue line denotes the KSS bound and the dashed red line denoted the fitting function $4\pi(\eta^{\prime}/s^{\prime})=e^{-0.33}\times(CT)^{-1.03}$. The right panel shows the same quantities in a log-log plot. The blue shaded area schematically represents the region where the entropy is negative, indicating the breakdown of the quantum-corrected $\eta/s$. We have set $\kappa=1$, $L=0.1$, $r_{0}=1$.

### v3

@@ 1 Introduction
Thus, our logic in this paper is the following. We compute quantum corrections to the IR retarded Green’s function $\mathcal{G}_{R}(\omega,T)$, working in two different temperature regimes: one which we refer to as semiclassical, for which the temperature scale dominates over the quantum scale, and one which we refer to as quantum, in which the quantum effects dominate. We should note that, despite being in a low temperature regime, we always work with $\omega\ll T$ to ensure that hydrodynamics is well defined. We then relate $\mathcal{G}_{R}(\omega,T)$ to the retarded Green’s function $G_{R}(\omega,T)$ in the UV, using Faulkner:2009wj ; Faulkner:2013bna ; Faulkner:2010da , and use Kubo’s formula to extract the shear viscosity.
We should stress that – unlike in the semiclassical regime, which is well-understood – our computation of $\eta$ in the deep quantum regime should be interpreted with caution, since a new framework is likely needed to accurately capture hydrodynamics. In this paper we will not address this challenge, but rather present our computations in both regimes, hoping that the results in the quantum regime might also provide guidance.
Moreover, we should note that it’s possible that certain instabilities will keep one from even accessing the extremal regime Buchel:2025ves ; Buchel:2025jup ; Gladden:2024ssb ; Choi:2024xnv , invalidating the need to incorporate the quantum fluctuations of the near-extremal brane. However, our goal in this paper is to investigate the quantum corrections of $\eta/s$ that arise under the assumption that the near-extremal AdS2 throat is stable. Provided this is the case, our analysis characterizes the universal IR physics associated with JT gravity, which is independent of the detailed UV completion. Finally, for completeness one should examine when the quantum effects computed here start competing with those due to higher-derivative operators in the gravitational theory (see Buchel:2026xtn for an efficient framework for extracting the behavior of e.g. $\eta/s$ in such theories). We will not examine these questions further in this paper.
@@
As we will see, the quantum corrections will generate a non-trivial temperature dependence for $\eta/s$, unlike the tree-level case for which $\eta/s=1/4\pi$. Moreover, we will find that $\eta/s$ reaches a minimum which violates the KSS bound in the semi-classical regime (and not in the deep quantum regime), suggesting that even a small amount of quantum corrections might lead to a lower bound. We will show that the minimum can be traced to a maximum in the quantum corrected entropy, while the viscosity itself is monotonic. We should also mention that recent studies PandoZayas:2025snm ; Nian:2025oei ; Gouteraux:2025exs ; Kanargias:2025vul have investigated quantum corrections to hydrodynamic transport coefficients arising from the JT mode at low temperatures [^4] The earlier work Jensen:2016pah also attempted to rewrite a generic theory of gravity near the AdS2 throat as a novel hydrodynamics coupled to the correlation functions of a conformal quantum mechanics.. As we show explicitly in Section 4, our results for $\eta/s$ are in agreement with those of PandoZayas:2025snm ; Kanargias:2025vul . The analysis of Gouteraux:2025exs considered a different conformal dimension from the one studied here and in PandoZayas:2025snm ; Kanargias:2025vul , and therefore yields different results.
@@ 4.2 Quantum corrected retarded Green’s function
In what follows, we will compute the Wightman Green’s function $\langle\mathcal{G}(\omega)\rangle$ in two distinct temperature regimes, and use the KMS condition
${\rm Im}\;\langle\mathcal{G}_{R}(\omega,\beta)\rangle=\frac{1}{2}\left(1-e^{-\beta\omega}\right)\langle\mathcal{G}(\omega)\rangle\,$, (4.13)
to obtain the corresponding retarded Green’s functions. First, we will examine the semi-classical limit $T\gg\frac{1}{C}$ in which the quantum effects are much smaller than the temperature scale, and then the opposite quantum regime $T\ll\frac{1}{C}$ in which quantum effects dominate over temperature. However, one should keep in mind that the results obtained in the quantum regime should be interpreted with caution, as a new theoretical framework may be needed for a proper description of hydrodynamics and of the shear viscosity (we return to this point in the Conclusions). Nonetheless, we present our $\eta/s$ results for $TC\ll 1$ since they may still offer useful physical intuition, with the understanding that a more thorough analysis could lead to a different behavior in this particular temperature range. Throughout the analysis we focus on the regime $\omega\ll M$, which implies [^11] One can show that $M\propto T^{2}$ and $M\propto T$ at the saddle points, in the semiclassical and quantum regimes respectively. $\omega\ll T$, ensuring that the frequency remains the lowest energy scale in the system and that hydrodynamics is well-defined (subject to the caveat above).
@@
    We evaluate this integral using the saddle-point approximation and a Fourier transformation to obtain the quantum-corrected retarded Green’s function
    $\hskip-28.45274pt\langle\mathcal{G}_{R}(\omega,\beta)\rangle$ $=i\frac{1}{2}\left(1-e^{-\beta\omega}\right)$
    $\times\frac{(2\pi)^{2\ell-1}}{{\rm sin}\left(\pi\ell+\pi i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}\textrm{exp}\left[\frac{\omega}{2}\times 2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]$
    $\times\left[2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]^{1-2\ell}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}{\Gamma\left(1-\ell-i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}\,.$ (4.18)
@@
    We evaluate this integral using the saddle-point approximation and performing a Fourier transformation, after which we obtain the following quantum-corrected retarded Green’s function
    $\langle\mathcal{G}_{R}(\omega,\beta)\rangle$ $=i\frac{1}{2}\left(1-e^{-\beta\omega}\right)\frac{(\ell-1)!^{2}}{2^{-5/2}e\pi}\frac{(2C)^{1-2\ell}}{\Gamma(2\ell)}\left(\frac{\beta}{2C}\right)^{-\frac{1}{2}}$
    $\times e^{\frac{\omega}{2}\pi\sqrt{2C\beta}}\frac{\pi}{{\rm sin}(\pi\ell+\pi i\frac{\omega}{2}\sqrt{2C\beta})}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}\pi\sqrt{2C\beta}\right)}{\Gamma(1-\ell-i\frac{\omega}{2}\sqrt{2C\beta})}\,.$ (4.21)
    Details of the derivation are described in Appendix A.2. Also, in Appendix A.3 we discuss a subtlety that arises from the fact that in this limit the Wightman Green’s function displays a period that is not set by $\beta$, but rather $\pi\sqrt{2C\beta}$. This in turn has implications for the KMS condition.
@@ 4.2.1 A quantum corrected $\ell^{\prime}$
    $\hskip-28.45274pt\langle\mathcal{G}_{R}(\omega,\beta)\rangle$ $=i\frac{1}{2}\left(1-e^{-\beta\omega}\right)$
    $\times\frac{(2\pi)^{2\ell-1}}{{\rm sin}\left(\pi\ell+\pi i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}\textrm{exp}\left[\frac{\omega}{2}\times 2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]$
    $\times\left[2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]^{1-2\ell}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}{\Gamma\left(1-\ell-i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}\,.$ (4.23)
@@
    $\hskip-28.45274ptF_{1}(\beta,\ell,\omega)$ $=-\frac{\Gamma(\frac{3}{2}-\ell)}{\Gamma(\frac{1}{2}+\ell)}(\pi)^{2\ell-1}(2C)^{1-2\ell}\frac{\Gamma(\ell-\frac{i\omega}{2\pi T})}{\Gamma(1-\ell-\frac{i\omega}{2\pi T})}\,$, (4.25)
    $\hskip-28.45274ptF_{2}(\beta,\ell,\omega)$ $=i\frac{1}{2}\left(1-e^{-\beta\omega}\right)\frac{(2\pi)^{2\ell-1}(2C)^{1-2\ell}}{{\rm sin}\left(\pi\ell+\pi i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}$
    $\times\textrm{exp}\left[\frac{\omega}{2}\times 2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}{\Gamma\left(1-\ell-i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}\,.$ (4.26)
@@
    $\ell_{1}^{\prime}=\frac{1}{2}-\frac{{\rm log}\left(\frac{F_{2}(\beta,\ell,\omega)}{F_{1}(\beta,\ell,\omega)}\right)+(1-2\ell)\,{\rm log}\left[\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]}{2\,{\rm log}\,\frac{\beta}{2C}}\,$, (4.28)
@@
    $\langle\mathcal{G}_{R}(\omega,\beta)\rangle$ $=i\frac{1}{2}\left(1-e^{-\beta\omega}\right)\frac{(\ell-1)!^{2}}{2^{-5/2}e\pi}\frac{(2C)^{1-2\ell}}{\Gamma(2\ell)}\left(\frac{\beta}{2C}\right)^{-\frac{1}{2}}$
    $\times e^{\frac{\omega}{2}\pi\sqrt{2C\beta}}\frac{\pi}{{\rm sin}(\pi\ell+\pi i\frac{\omega}{2}\sqrt{2C\beta})}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}\pi\sqrt{2C\beta}\right)}{\Gamma(1-\ell-i\frac{\omega}{2}\sqrt{2C\beta})}\,.$ (4.29)
@@
    $\langle\mathcal{G}_{R}(\omega,\beta)\rangle$ $=F_{3}(\beta,\ell,\omega)\times\left(\frac{\beta}{2C}\right)^{-\frac{1}{2}}=F_{1}(\beta,\ell,\omega)\times\left(\frac{\beta}{2C}\right)^{1-2\ell_{2}^{\prime}}\,$, (4.30)
@@
    $\hskip-28.45274ptF_{3}(\beta,\ell,\omega)$ $=i\frac{1}{2}\left(1-e^{-\beta\omega}\right)\frac{(\ell-1)!^{2}}{2^{-5/2}e\pi}\frac{(2C)^{1-2\ell}}{\Gamma(2\ell)}e^{\frac{\omega}{2}\pi\sqrt{2C\beta}}\frac{\pi}{{\rm sin}(\pi\ell+\pi i\frac{\omega}{2}\sqrt{2C\beta})}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}\pi\sqrt{2C\beta}\right)}{\Gamma(1-\ell-i\frac{\omega}{2}\sqrt{2C\beta})}\,.$ (4.31)
@@
    $\ell_{2}^{\prime}=\frac{3}{4}+\frac{1}{2}\frac{\log{\left(\frac{F_{1}(\beta,\ell,\omega)}{F_{3}(\beta,\ell,\omega)}\right)}}{\log{\left(\frac{\beta}{2C}\right)}}\,.$ (4.32)
@@ 4.3 Quantum corrected shear viscosity and entropy density
To better illustrate the behavior of the shear viscosity, we write down its approximated expression at high temperature $T\gg\frac{1}{C}$ for $\ell=1$,
@@
where $\cdots$ denotes $\mathcal{O}(r_{0}^{-2})$ and $\mathcal{O}((CT)^{-2})$ terms. We note that this expression (4.37) for the quantum-corrected $\eta^{\prime}$ in the semiclassical region is consistent with the complementary work PandoZayas:2025snm ; Kanargias:2025vul . However, the full expression (4.3) is generically more involved, and captures additional effects to those that appear to leading order in the expansion (4.37).
On the other hand, in the quantum regime $T\ll\frac{1}{C}$, taking $\ell=1$ we have
$\eta^{\prime}=\frac{1}{2\kappa^{2}}\left(\frac{r_{h}}{L}\right)^{2}T^{2\ell_{2}^{\prime}-2}$ $=\frac{1}{2\kappa^{2}}\left(\frac{r_{h}}{L}\right)^{2}\times\frac{2^{3/2}}{e\pi}\frac{1}{\sqrt{2CT}}$ (4.38)
$\approx\frac{1}{2\kappa^{2}}\left(\frac{r_{0}}{L}\right)^{2}\times\frac{2^{3/2}}{e\pi}\frac{1}{\sqrt{2CT}}+\cdots\,.$ (4.39)
We will come back to these expressions for the shear viscosity in the next section, where we make a direct comparison to the behavior of the absorption cross-section. Also, we should stress that when we plot the viscosity to entropy ratio we will use the full expression (4.3) for the shear viscosity, and not the approximations above. In Figure 2 we plot the ratio of the quantum-corrected shear viscosity $\eta^{\prime}$ to the classical shear viscosity $\eta$ in the two distinct temperature regimes.
@@
Figure 5: The left panel shows the temperature dependence of the quantum-corrected shear viscosity-entropy ratio in the quantum regime $CT\ll 1$. The dashed blue line denotes the KSS bound and the dashed red line denoted the fitting function $4\pi(\eta^{\prime}/s^{\prime})\approx 0.24(CT)^{-0.5}$. The right panel shows the same quantities in a log-log plot. The blue shaded area schematically represents the region where the entropy is negative, indicating the breakdown of the quantum-corrected $\eta/s$. We have set $\kappa=1$, $L=0.1$, $r_{0}=1$.